Two waves are generated on a string of length 3.2What is the answer for (b) and (c) ??
Answer #1

for 3 loops,
L = 1.5*wavelength
3.2 = 1.5*wavelength
wavelength = 2.13 m

b)
k' = 2*pi/wavelength
= 2*pi/2.13
= 2.95 m-1

c)
w' = v*k'
= 111*2.95
= 326.9 rad/s